Grade 12th passAlgebra

Simplify the expression and find its value when a = 5 and b = – 3.
2(a2 + ab) + 3 – ab

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question it is given that a = 5 and b = -3
We have,
= 2a2+ 2ab + 3–ab
= 2a2 + ab + 3
Then, substitute the value of a and b in the equation
= (2 × 52) + (5 × (-3)) + 3
= (2 × 25) + (-15) + 3
= 50 - 15 + 3
= 53–15
= 38

Thanks